consider that we have a circuit with a potential difference of 3.2 volts with a current of 0.6 a ammeter the internal resistance of the battery at 0.5 homes use emf formula

1
Expert's answer
2022-01-14T09:48:05-0500

Determine the electromotive force:


"E=V+Ir=3.2+0.6·0.5=3.5\text{ V}."